10 gram of ice at -10 degree Celsius is added to 10 gram of water at 85 degree Celsius. What is the final temperature and amount of ice left in the system? (System is kept inside an ideal insulator).

Grade:10

1 Answers

Sanju
6 years ago
HEAT GAINED = HEAT LOST
use eqn mcdt....m is the mass [0.01 kg], c is the specific heat capacity of water [4200 J kg-1 K-1] and dt is change in temp.
T= -10 0C = 263K and T= 85 0C = 358K
Let T be the final temp.
Then,  0.01 kg* 4200 * (T-263) = 0.01 kg * 4200 * (358-T)    [Heat gained = heat lost]
Solving the eqn gives
T-263 = 358-T
2T = 358+263 = 621
or, T = 310.5K = 37.35 0C
FINAL TEMP, T = 310.5K = 37.35 0C
NO ICE WILL LEFT IN THE SYSTEM AS IT COMPLETELY  BECOMES WATER